OpenStudy (anonymous):

f(x) {x+5 if -4 is less than or equal to x which is less than 1 {9 if x=1 {-x+4 if x>1 find the domain and intercepts

OpenStudy (anonymous):

@modphysnoob

OpenStudy (anonymous):

@bahrom7893 @Eyad @e.mccormick

OpenStudy (anonymous):

f(x) { x+5 x=[-4,1) {9 if x=1 {-x+4 if x>1 \[domain = [-4,+\infty )\] intercept x=4

OpenStudy (anonymous):

how did you find that?

OpenStudy (anonymous):

{x+5 if -4 is less than or equal to x which is less than 1 that means x =[-4,1) and we have the rest domains of x so the main domain is the sum all of them that is [−4,+∞)

OpenStudy (anonymous):

thank you :)